OpenStudy (anonymous):

Use elimination to solve the system. What is the value of y?: 3x+5y=16 and 8x-5y=28

OpenStudy (phi):

though indirect, I would add the 2 equations together. can you do that ?

OpenStudy (anonymous):

Ohh add. Yup!

OpenStudy (anonymous):

I was subtracting.>~>

OpenStudy (anonymous):

y=4

OpenStudy (phi):

it is easy to add. put a + sign in, like this 3x+5y=16 8x-5y=28 3x+8x +5y + -5y = 16+28 that simplifies to 11x=44 x=4 note you want y , not x

OpenStudy (anonymous):

Okay. I tend to lock up with math. Thanks!:D

OpenStudy (phi):

what do you get for y ?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Er sorry!

OpenStudy (phi):

I hope you notice x=4. y is not 4

OpenStudy (jhannybean):

hint: plug x=4 into any one of the two equations to solve for y.

OpenStudy (anonymous):

Ohmygod. Obviously. So sorry.._.

OpenStudy (jhannybean):

3(4) +5y = 16 5y = 16 - 3(4) y = (16-3(4))/5

OpenStudy (jhannybean):

or plug x=4 into the other equation. 8(4) -5y = 28 -5y = 28 -(8(4)) y= [28-8(4)] / (-5)

OpenStudy (anonymous):

Okay got it. Thank you!c: